Please answer any part
Question: 1 pt 1 of 10 (6 complete) This Quiz: 10 pts possibk Avwwty ofpropon?are berg ideed by man nest to anagement action and stock valu· RE:H Corporators most recent didend was S1 64 per share bs expected annual rate of dded gowthis5%?andthe repred renumts now 15% direct the s avwities. Determine the impadt on share price for each of the folowing proposed actions . Do nolhing which will leave the key financial variables unchanged Invest r, a new machine Bui wll naease dvdend growth rate to 7% and lower the requred retum to 12% Merge wth another firm which wil reduce the growth tate to 3% and tame Po roared return to 16% If the tem does noting thait w leave the key nancial vanables unchanged, the value of the fim will beRound to the mearest cent) the t-mests n a new madre tat wo ncrease me dedend own rate to 7% and Ower he reared rehm to 12%. teyn, se nrm we besC] ?Remote sent on
Answer #1

Ans a) Value of share = dividend in next year/ ( required rate - growth rate)

= 1.64 * ( 1+ .05) /(.15 - .05)

= $17.22

Since the number of stock is not given it will be difficult to find value of firm but it will be equal to share price* value of share.

Ans b) Value of share = 1.64 * 1.07 /(.12 - .07)

= $ 35.1

Ans c) Value of share = 1.64 * 1.09/(.16 - .09)

= $25.54

Ans d) Value of share = 1.64 * 1.03 /(.16 - .03)

= $ 12.99

ans e) Value of share = 1.64 * 1.08 /(.16 - .08)

= $ 22.14